# Accelerating PoT Quantization on Edge Devices

### v1

@@ Abstract
In general, a single bit-shift operation is computationally cheaper than a multiplication, but different PoT quantization methods introduce varying numbers of shifts and additions to replace a single multiplication. For example, the methods from Przewlocka-Rus et al. [4] and QKeras [5] use a single shift operation to replace a multiplication. Other methods like APoT [2] and MSQ [3] replace a multiplication by two shift operations and one addition. Such variations in PoT quantization methods introduce different complexities in hardware design, including the processing elements (PEs).
While previous works on PoT quantization [2, 3, 4] focused on achieving accuracy comparable to floating-point, they did not provide an analysis of their hardware efficiency. Therefore, there is a lack of discussion on hardware design implementations of different PoT quantization methods on edge devices. Furthermore, another important gap we identified is the lack of open-source pipelines to explore different PoT quantization methods for DNNs, accelerate them using custom hardware and evaluate them on resource-constrained edge devices.
@@ II-A Power-of-Two Quantization
Previous works proposed different PoT quantization methods for DNNs to show efficacy over uniform quantization. We group them based on the number of PoT terms to represent each quantization level. The first group ( [7, 8, 9, 4, 10]) uses a single PoT term per quantization level, and the second ( [11, 12, 2, 13, 14, 3, 15, 16]) can have multiple PoT terms per quantization level. In general, the shift terms derived from the PoT terms are represented within to bits. A key challenge is to understand how the bit-width and the number of PoT terms allocated per quantization level affect the design of an associated hardware accelerator, and more specifically the PE design. Previous works do not provide such analysis.
In this paper, we aim to address this key point. We opted for 4-bit as the bit-width, since it is a commonly used configuration for low-bit PoT quantization methods [4]. In addition, while previous works investigated joint activation and weight quantization [11, 15, 16], we consider 4-bit PoT quantization for weights and 8-bit uniform quantization for activations; this avoids the additional overhead of applying PoT quantization to the activations during inference.
@@ II-B Processing Element Design for PoT Quantization
The core computational unit of a DNN accelerator is the PE that executes the multiply-accumulate operations (we call it mult-PE). Hence, optimizing the mult-PE design can enhance the overall efficiency of the accelerator. Shift-based accelerators for PoT quantization replace mult-PEs with shift-PEs. For example, JLQ [10] proposed a shift-PE design based on their jumping log quantization. While the design is optimized, the circuit is inefficient in handling activations based on two’s complement. Liu et al. [13] proposed a BS-MUL architecture for 16-bit activations and 12-bit weights. However, their target bit-width differs from ours (4-bit) and their circuit exhibits inefficiencies in handling negative activations. Vogel et al. [12] proposed a shift-PE based on two 8-bit shift terms, hence the accumulation in the PE circuit was expensive. Przewlocka-rus et al. [4] discussed the efficiency of shift-PEs based on single and double shift terms, but without providing a hardware design for them. In summary, none of the aforementioned works compare the hardware design for different PoT quantization methods. We address this in Section III.

### v2

@@ Abstract
In general, a single bit-shift operation is computationally cheaper than a multiplication, but different PoT quantization methods introduce varying numbers of shifts and additions to replace a single multiplication. For example, the methods from Przewlocka-Rus et al. [4] and QKeras [5] use a single shift operation to replace a multiplication. Other methods like APoT [2] and MSQ [3] replace a multiplication by two shift operations and one addition. Such variations in PoT quantization methods introduce different complexities in hardware design, including the processing elements (PEs). Therefore, while previous works on PoT quantization [2, 3, 4] focused on achieving accuracy comparable to floating-point, they provide limited discussion on hardware design complexity of the PEs for the different PoT quantization methods on edge devices. Furthermore, another important gap we identified is the lack of open-source pipelines to explore different PoT quantization methods for DNNs, accelerate them using custom hardware and evaluate them on resource-constrained edge devices.
@@ II-A Power-of-Two Quantization
Previous works proposed different PoT quantization methods for DNNs to show efficacy over uniform quantization. Some works ( [7, 8, 9, 4, 10]) use a single PoT term per quantization level, while others ( [11, 12, 2, 13, 14, 3, 15, 16]) can have multiple PoT terms per quantization level. In general, the shift terms derived from the PoT terms are represented within to bits. A key challenge is to understand how the bit-width and the number of PoT terms allocated per quantization level affect the design of an associated hardware accelerator, and more specifically the PE design. Previous works do not provide such analysis.
In this paper, we address this key point. We opted for 4-bit PoT quantization for weights and 8-bit uniform quantization for activations, since it is a commonly used configuration for low-bit PoT quantization methods [4]; this avoids the additional overhead of applying PoT quantization to the activations during inference. Note that previous works investigated joint activations and weights quantization [11, 15, 16].
@@ II-B Processing Element Design for PoT Quantization
The core computational unit of a DNN accelerator is the PE that executes the multiply-accumulate operations (we call it mult-PE). Hence, optimizing the mult-PE design can enhance the overall efficiency of the accelerator. Shift-based accelerators for PoT quantization replace mult-PEs with shift-PEs. For example, JLQ [10] proposed a shift-PE design based on their jumping log quantization. While the design is optimized, the circuit is inefficient in handling activations based on two’s complement. Liu et al. [13] proposed a BS-MUL architecture for 16-bit activations and 12-bit weights. However, their target bit-width differs from ours (4-bit) and their circuit exhibits inefficiencies in handling negative activations. Vogel et al. [12] proposed a shift-PE based on two 8-bit shift terms, hence the accumulation in the PE circuit was expensive. While none of the aforementioned works compare the hardware design of shift-PEs for different PoT quantization methods, Przewlocka-Rus et al. ( [4, 17]) discussed the efficiency of shift-PEs based on single and double shift terms and presented a hardware design for single shift-term PE, but did not provide a design for double shift-term PE. In our work, we provide hardware design and efficiency analysis of single and double shift-term PEs for different PoT quantization methods in Section III.